A forty foot ladder is set against the wall. the distance of the base of the ladder to the wall is equal to the height of the ladder. How high is the ladder on the wall?
If I was to use the theorem, my answer is 0. (40)^2-(40)^2=b^2 1600-1600=b^2
0=b^2.
On the drawing it shows a 90o angle between the ladder and the wall.

Answer by mananth(16946)

let distance of foot of ladder be x
height will also be x
ladder length = 40
The ladder ,ground and the wall form a right triangle
ground distance = leg2
height at which ladder touches
Hypotenuse = ladder length
Pythagoras theorem states that

Hypotenuse ^2= leg1^2+leg2^2
40^2=x^2+x^2
1600=2x^2
1600/2= x^2
800=x^2
take the square root
x=28.285 feet